OpenStudy (anonymous):

what is 18 1/15 - 9 3/5

OpenStudy (anonymous):

you mean 18/15 - 27/5? if so it equals 4.2 or 21/5

OpenStudy (yuki):

is it\[18 {1 \over 15} - 9 {3 \over 5}\] ?

OpenStudy (anonymous):

yes yuki

OpenStudy (yuki):

mixed numbers are fairly easy when you understand that it can also be changed into an equivalent improper fraction

OpenStudy (anonymous):

ooh 271/15 - 48/5 = 31/5 ? lol i think

OpenStudy (anonymous):

which is 6 5/1

OpenStudy (yuki):

for example, \[18 {1 \over 15}= 18 + {1 \over 15}\] so \[{18*15+1 \over 15} = {217 \over 15}\]

OpenStudy (yuki):

you do the same thing with the other fraction, but before we do that, we need to make the denominator common. so instead of \[9 {3 \over 5}\] we will use \[9 {9 \over 15}\] which as an improper fraction, is \[{9*15+9 \over 15} = {144 \over 15}\]

OpenStudy (anonymous):

Who elsee gott daht?

OpenStudy (yuki):

now if you subtract them, \[{217 - 144 \over 15} = {73 \over 15}\]

OpenStudy (yuki):

if you need it as a mixed number, 73 / 15 = 4 remainder 13 so \[4 {13 \over 15}\]

OpenStudy (anonymous):

I GOT DIFF.